# K theoretic virtual fundamental cycle of global Kuranishi chart

### v1

@@ 3.4. Push-forward from inertia orbifold to point
Restricting or extending a symbol outside its support preserves the analytic index by the excision property (Theorem 3.7 in [3]). So taking into account Proposition 2.1, we only need to prove the proposition when $\mathcal{Y}=\mathcal{W}$ is closed. Given $u\in K^{0}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})$, let $u_{\mathbf{g}}=e_{\mathbf{g}}^{-1}i_{\mathbf{g}}^{\ast}(str_{g}u)$. With the assumption that $\mathcal{Y}$ is closed, we compute $\operatorname{ind}u_{+}$ by Kawasaki’s index theorem:
$\operatorname{ind}u_{+}=\sum_{(g)}(-1)^{\dim\mathcal{Y}^{(g)}}\int_{[T\mathcal{Y}^{(g)}]}ch(u_{g})_{+}\cdot td(T\mathcal{Y}^{(g)}\otimes\mathbb{C}).$
Note that supertraces act as identity on $u_{g}$.
Each term on the right-hand side is (up to a sign $(-1)^{\dim\mathcal{Y}^{(g)}}$) the principle term (that corresponds to the non-twisted strata $g_{1}\in\left<g\right>$, the group generated by $g$) of Kawasaki’s index formula for $u_{g}$ over $\mathcal{Y}^{(g)}$.
$\operatorname{ind}[u_{g}]_{+}=\sum_{(g_{1})}(-1)^{\dim\mathcal{Y}^{(g)(g_{1})}}\int_{[T\mathcal{Y}^{(g)(g_{1})}]}ch\left(\frac{i_{g_{1}}^{\ast}u_{g}}{str_{g_{1}}\wedge^{\ast}\mathcal{N}^{\ast}_{g_{1}}\otimes\mathbb{C}}\right)_{+}\cdot td(T\mathcal{Y}^{(g)(g_{1})}\otimes\mathbb{C}).$
Note that all terms with $\mathcal{Y}^{(g)(g_{1})}=\mathcal{Y}^{(g)}$ in the sum are the same. We move terms corresponding to $g_{1}\notin\left<g\right>$ to the left and divide both sides by $m(g)$ to get
$\begin{array}[]{ll}&\displaystyle\int_{[T\mathcal{Y}^{(g)(g_{1})}]}ch\left(\dfrac{i_{g_{1}}^{\ast}u_{g}}{str_{g_{1}}\wedge^{\ast}\mathcal{N}^{\ast}_{g_{1}}\otimes\mathbb{C}}\right)_{+}\cdot td(T\mathcal{Y}^{(g)(g_{1})}\otimes\mathbb{C})\\ =\dfrac{1}{m(g)}\operatorname{ind}[u_{g}]_{+}-&\displaystyle\sum_{(g,g_{1})\in\mathcal{G}}(-1)^{\dim\mathcal{Y}^{(g)(g_{1})}}\displaystyle\int_{[T\mathcal{Y}^{(g)(g_{1})}]}ch\left(\dfrac{i_{g_{1}}^{\ast}u_{g}}{str_{g_{1}}\wedge^{\ast}\mathcal{N}^{\ast}_{g_{1}}\otimes\mathbb{C}}\right)_{+}\cdot td(T\mathcal{Y}^{(g)(g_{1})}\otimes\mathbb{C}).\end{array}$
This expresses the desired integration on the sector $\mathcal{Y}^{(g)}$ in terms of the index of $u_{g}$ and integrations of the same form on $2$-sectors. The same result with $(\mathcal{Y}^{(\mathbf{g})},\mathcal{Y}^{(\mathbf{g})(g_{n+1})},u_{\mathbf{g}})$ in the place of $(\mathcal{Y}^{(g)},\mathcal{Y}^{(g)(g_{1})},u_{g})$ expresses relevant integrations on $(n+1)$-sectors in terms of $\operatorname{ind}[u_{\mathbf{g}}]_{+}$ and integrations on $(n+2)$-sectors. Since points in the compact orbifold $\mathcal{Y}$ have finite isotropy, there is a uniform bound $N$ on the order of the isotropy groups at points in $\mathcal{Y}$, and $n$-sectors $\mathcal{Y}^{\mathbf{g}}$ become empty when $n>N$. Thus, we can write
$ind([u]_{+})=\sum_{\mathbf{g}\in\mathcal{G}}\frac{(-1)^{|\mathbf{g}|+\dim\mathcal{Y}^{(g)}}}{m(\mathbf{g})}ind([u_{\mathbf{g}}]_{+})$
as a finite sum.
@@
This likewise follows from an induction on $|\mathbf{g}|$, where the base case and the induction step are both
@@
for appropriate $\mathcal{N},u$. Indeed, the latter is equivalent to
$\pi_{g}^{\ast}i_{g}^{\ast}\left(str_{g}\frac{u}{\wedge^{\ast}(\mathcal{N}_{g}^{\ast}\otimes\mathbb{C})}\right)\cdot str_{g}\wedge^{\ast}(\mathcal{N}_{g}^{\ast}\otimes\mathbb{C})=str_{g}u,$
which is
$\left(\pi_{g}^{\ast}i_{g}^{\ast}\bigg{(}str_{g}\frac{u}{\wedge^{\ast}(\mathcal{N}_{g}^{\ast}\otimes\mathbb{C})}\bigg{)}-\bigg{(}str_{g}\frac{u}{\wedge^{\ast}(\mathcal{N}_{g}^{\ast}\otimes\mathbb{C})}\bigg{)}\right)\cdot str_{g}\wedge^{\ast}(\mathcal{N}_{g}^{\ast}\otimes\mathbb{C})=0.$
The first term vanishes on $\mathcal{Y}^{g}$, and the second term vanishes on $\mathcal{N}^{g}\setminus\mathcal{Y}^{g}$.
Finally, applying part (1) of Proposition 3.3, we know that
$\begin{array}[]{ll}ind([u_{\mathbf{g}}]_{+})&=p_{\ast}^{fake}(\mu(u_{\mathbf{g}}))\\ &=p_{\ast}^{fake}i_{\mathbf{g}}^{!}\mu(i_{\mathbf{g}})_{!}(u_{\mathbf{g}})\\ &=p_{\ast}^{fake}(i_{\mathbf{g}}^{!}(\mu(u))^{(\mathbf{g})}),\end{array}$
@@ 3.5. Compatibility
Given an element $\alpha\in K^{0}(\mathcal{Y})$ and an element $T\in K_{0}(\mathcal{Y})$, we define their pairing
@@
The following proposition follows and Propositions 3.2 and 3.4.
@@ 4. K-theoretic virtual fundamental cycle
A stably almost complex global Kuranishi chart consists of a quadruple $A=(G,Y,E,s)$, where $G$ is a compact Lie group, $Y$ a locally $C^{1}$ $G$-manifold with almost free $G$-action, such that $T^{\ast}_{G}Y$ is stably complex, $E$ a locally $C^{1}$ $G$-equivariant stably complex vector bundle over $Y$, and $s$ an equivariant section $Y\to E$.
In [2], Abouzaid, Mclean and Smith constructed $stably$ almost complex global Kuranishi charts on the compactified moduli space of stable maps ¯$\bar{\mathcal{M}}_{g.n}(X,d)$ to general symplectic targets $X$. We shall make the following modifications and use orbifold Kuranishi charts.
@@
An almost complex global orbifold Kuranishi chart consists of a triple $\mathcal{A}=(\mathcal{Y,E},\sigma)$, where $\mathcal{Y}$ a smooth, compact orbifold with boundary, $\mathcal{E}$ a complex smooth orbi-bundle over $\mathcal{Y}$, and $\sigma$ a continuous section $\mathcal{Y}\to\mathcal{E}$ that vanishes nowhere on $\partial\mathcal{Y}$.
Remark that almost complex global orbifold Kuranishi charts exist. Since the real dimensions of $Y$ and fibers of $E$ have the same parity, we can always add a same $\mathbb{R}^{k}$ to them (and identity map $id^{k}:\mathbb{R}^{k}\to\mathbb{R}^{k}$ to the section $s$) to get complex $Y,E$’s. An almost complex global Kuranishi chart $(G,Y,E,s)$ gives an almost complex orbifold Kuranishi chart $(\mathcal{Y,E},\sigma)$ by quotient out the $G$ action and shrinking to a compact neighborhood of $\sigma^{-1}(0)$. $C^{1}$ structures induce smooth structures, so we can also modify $Y,E$ to smooth orbifolds and smooth bundles. Note that $s$ is not smooth in general.
By [2] Proposition 4.62 and Lemma 3.4 (see also [1]), global Kuranishi charts $A=(G,Y,E,s)$ for the same moduli space of stable maps of a certain type are connected by a sequence of operations of the following types, or their inverses.
@@
    Free quotients $A\to A/H$, with $H$ a normal subgroup of $G$. This replaces $(G,Y,E,s)$ by $(G/H,Y/H,E/H,s)$, where $E/H$ is viewed as a $G/H$-bundle on $Y/H$.
@@
    The induced map $A\to A\times_{G}G^{\prime}$ associated to an embedding $G\to G^{\prime}$. This replaces $(G,Y,E,s)$ by $(G^{\prime},Y\times_{G}G^{\prime},\pi^{\ast}E,\pi^{\ast}s)$, where $\pi:Y\times_{G}G^{\prime}\to Y$ is the projection induced by a slicing $G^{\prime}\to G$.
@@
    The stabilization by a $G$-equivariant vector bundle $\pi:W\to Y$, replaces $(G,Y,E,s)$ by $(G,W,\pi^{\ast}E\oplus\pi^{\ast}W,\pi^{\ast}s\oplus\Delta)$, where $\Delta$ is the tautological ‘diagonal’ section of $W\to\pi^{\ast}W$.
@@
    An open embedding $A^{\prime}\to A$ of a neighborhood of $s^{-1}(0)$. This replaces $(G,Y^{\prime},E|_{Y^{\prime}},s|_{Y^{\prime}})$ by $(G,Y,E,s)$ for $Y^{\prime}\subseteq Y$ open containing $s^{-1}(0)$.
In turn, the orbifold Kuranishi charts are related by a sequence of the following operation and its inverse.
@@
We define K-theoretic virtual fundamental cycle $\mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}\in K_{0}(\mathcal{Y})$ of an almost complex global orbifold Kuranishi chart $(\mathcal{Y,E},\sigma)$ as follows. Let $\pi:T^{\ast}\mathcal{Y}\to\mathcal{Y}$ be the projection that forgets the cotangent vector. Let $\mathcal{S}=\wedge^{\ast}_{\mathbb{C}}T^{\ast}\mathcal{Y},\theta=\wedge^{\ast}_{\mathbb{C}}\mathcal{E}$ and $\mathcal{O}=\mathcal{S}\otimes\theta$. Remark that $\mathcal{S},\theta$ are $\mathbb{Z}_{2}$-graded bundles and so is $\mathcal{E}$. We add Hermitian structures to $\mathcal{Y,E}$, and let
$\begin{array}[]{ll}p&=\lambda_{\xi}\otimes 1+\varepsilon 1\otimes\pi^{\ast}\lambda_{\sigma}:\pi^{\ast}\mathcal{O}\to\pi^{\ast}\mathcal{O},\\ \mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}&=\mu(p)\end{array}.$
Recall that $\varepsilon$ is the operator on $\mathbb{Z}_{2}$ graded objects acting as $(-1)^{a}$ on the direct summand of degree $a$. Also recall that $\lambda$ is as defined in the introduction, namely given a section $\tau:\mathcal{X}\to\mathcal{F}$ of a Hermitian bundle, we define $\lambda_{\tau}:\wedge^{\ast}\mathcal{F}\to\wedge^{\ast}\mathcal{F}$ by $\tau(x)\wedge\cdot+i_{\tau(x)}(\cdot)$ over points $x\in\mathcal{X}$, where $i_{\tau(x)}$ is the contraction with $\tau(x)$ via the Hermitian metric on $\mathcal{F}$. Form $\lambda_{\xi}:\pi^{\ast}\mathcal{S}\to\pi^{\ast}\mathcal{S}$ and $\lambda_{\sigma}:\mathcal{E}\to\mathcal{E}$ corresponding to the diagonal map $\xi:T^{\ast}\mathcal{Y}\to\pi^{\ast}T^{\ast}\mathcal{Y}$ and $\sigma:\mathcal{Y}\to\mathcal{E}$. Observe that $\lambda_{\xi}$ (resp. $\lambda_{\sigma}$) are of degree $1$, so $p$ is also of degree $1$. Observe that $\lambda_{\xi},\lambda_{\sigma}$ are self-dual and squares to $|\xi|^{2}\cdot 1,|\sigma(x)|^{2}\cdot 1$, so $p^{\ast}p=(|\xi|^{2}+|\sigma(x)|^{2})\cdot 1$. Hence $p$ is invertible outside the compact set =$\mathcal{M}=\{|\sigma(x)|=0,|\xi|=0\}$ and therefore represents an element in $K^{0}(T^{\ast}\mathcal{Y})$.
It follows from Proposition 3.3 that in terms of the K-theoretic virtual fundamental cycle, the operation (*) corresponds to the push forward by $i_{\ast}$ (defined in Section 2.3) and then pull back by $j^{!}$ (defined in Section 3.3).
@@
If $\sigma$ is transversal to the zero section 00, then $\mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}=i_{\ast}[\sigma^{-1}(0)]$.
@@
Note that the space $\mathcal{M}=\sigma^{-1}(0)$ is stably almost complex. We can define its K-theoretic fundamental cycle by Bott periodicity theorem $K_{0}(\Sigma^{k}\mathcal{M})=K_{0}(\mathcal{M})$. This identification is exactly our $i^{!}$ for $i:\mathcal{M}\to B(\mathcal{M}\times\mathbb{R}^{2k})$. Next, by part (2) of Proposition 3.3, the push-forwards $i_{\ast}$ and the wrong-way pull-back $i^{!}$ are intertwined by the choose-an-operator map, thus proving the theorem. ∎
@@
$\mathcal{O}^{vir}_{\sigma},\mathcal{V}\right>=\operatorname{ind}[\mathcal{E}\otimes\pi^{\ast}\mathcal{V}]_{+}.$

### v2

@@ 3.5. Push-forward from inertia orbifold to point
Restricting or extending a symbol outside its support preserves the analytic index by the excision property (Theorem 3.7 in [3]). Taking into account Proposition 2.1, it suffices to prove the proposition when $\mathcal{Y}=\mathcal{W}$ is closed.
Given $u\in K^{0}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})$, define
$u_{\mathbf{g}}=e_{\mathbf{g}}^{-1}\,i_{\mathbf{g}}^{\ast}(str_{g}\,u).$
We compute $\operatorname{ind}u_{+}$ by Kawasaki’s index theorem:
$\operatorname{ind}u_{+}=\sum_{(g)}(-1)^{\dim\mathcal{Y}^{(g)}}\int_{[T\mathcal{Y}^{(g)}]}ch(u_{g})_{+}\cdot td\Bigl{(}T\mathcal{Y}^{(g)}\otimes\mathbb{C}\Bigr{)}.$
Note that $u_{g}$ is already a supertrace, so the supertraces act as the identity on $u_{g}$.
Each term on the right-hand side is (up to the sign $(-1)^{\dim\mathcal{Y}^{(g)}}$) the principal term (corresponding to the non-twisted stratum $g_{1}\in\langle g\rangle$, the group generated by $g$) of Kawasaki’s index formula for $u_{g}$ over $\mathcal{Y}^{(g)}$. By Kawasaki’s index formula,
$\operatorname{ind}[u_{g}]_{+}=\sum_{(g_{1})}(-1)^{\dim\mathcal{Y}^{(g)(g_{1})}}\int_{[T\mathcal{Y}^{(g)(g_{1})}]}ch\!\left(\frac{i_{g_{1}}^{\ast}u_{g}}{str_{g_{1}}\,\wedge^{\ast}(\mathcal{N}^{\ast}_{g_{1}}\otimes\mathbb{C})}\right)_{+}\cdot td\Bigl{(}T\mathcal{Y}^{(g)(g_{1})}\otimes\mathbb{C}\Bigr{)}.$
Note that all terms with $\mathcal{Y}^{(g)(g_{1})}=\mathcal{Y}^{(g)}$ in the sum are the same. We move the terms corresponding to $g_{1}\notin\langle g\rangle$ to the left and divide both sides by $m(g)$ to obtain
$\begin{array}[]{ll}&\displaystyle\int_{[T\mathcal{Y}^{(g)(g_{1})}]}ch\!\left(\dfrac{i_{g_{1}}^{\ast}u_{g}}{str_{g_{1}}\,\wedge^{\ast}(\mathcal{N}^{\ast}_{g_{1}}\otimes\mathbb{C})}\right)_{+}\cdot td\Bigl{(}T\mathcal{Y}^{(g)(g_{1})}\otimes\mathbb{C}\Bigr{)}\\[2.84526pt] =\dfrac{1}{m(g)}\operatorname{ind}[u_{g}]_{+}-&\displaystyle\sum_{(g,g_{1})\in\mathcal{G}}(-1)^{\dim\mathcal{Y}^{(g)(g_{1})}}\displaystyle\int_{[T\mathcal{Y}^{(g)(g_{1})}]}ch\!\left(\dfrac{i_{g_{1}}^{\ast}u_{g}}{str_{g_{1}}\,\wedge^{\ast}(\mathcal{N}^{\ast}_{g_{1}}\otimes\mathbb{C})}\right)_{+}\cdot td\Bigl{(}T\mathcal{Y}^{(g)(g_{1})}\otimes\mathbb{C}\Bigr{)}.\end{array}$
This equation expresses the desired integration on the sector $\mathcal{Y}^{(g)}$ in terms of the index of $u_{g}$ and integrations of the same form on 2-sectors. A similar result with $(\mathcal{Y}^{(\mathbf{g})},\,\mathcal{Y}^{(\mathbf{g})(g_{n+1})},\,u_{\mathbf{g}})$ in place of $(\mathcal{Y}^{(g)},\,\mathcal{Y}^{(g)(g_{1})},\,u_{g})$ expresses the relevant integrations on $(n+1)$-sectors in terms of $\operatorname{ind}[u_{\mathbf{g}}]_{+}$ and integrations on $(n+2)$-sectors.
@@
This also follows by induction on $|\mathbf{g}|$; the base case and the induction step both reduce to
@@
for appropriate $\mathcal{N}$ and $u$. Indeed, the latter is equivalent to
$\pi_{g}^{\ast}i_{g}^{\ast}\Bigl{(}str_{g}\,\frac{u}{\wedge^{\ast}(\mathcal{N}_{g}^{\ast}\otimes\mathbb{C})}\Bigr{)}\cdot str_{g}\,\wedge^{\ast}(\mathcal{N}_{g}^{\ast}\otimes\mathbb{C})=str_{g}\,u,$
which can be rewritten as
$\Bigl{(}\pi_{g}^{\ast}i_{g}^{\ast}\Bigl{(}str_{g}\,\frac{u}{\wedge^{\ast}(\mathcal{N}_{g}^{\ast}\otimes\mathbb{C})}\Bigr{)}-str_{g}\,\frac{u}{\wedge^{\ast}(\mathcal{N}_{g}^{\ast}\otimes\mathbb{C})}\Bigr{)}\cdot str_{g}\,\wedge^{\ast}(\mathcal{N}_{g}^{\ast}\otimes\mathbb{C})=0.$
Here, the first term vanishes on $\mathcal{Y}^{g}$ and the second term vanishes on $\mathcal{N}^{g}\setminus\mathcal{Y}^{g}$.
Finally, applying part (1) of Proposition 3.3, we obtain
$\begin{array}[]{ll}\operatorname{ind}([u_{\mathbf{g}}]_{+})&=p_{\ast}^{fake}\Bigl{(}\mu(u_{\mathbf{g}})\Bigr{)}\\[2.84526pt] &=p_{\ast}^{fake}\Bigl{(}i_{\mathbf{g}}^{!}\,\mu\bigl{(}(i_{\mathbf{g}})_{!}(u_{\mathbf{g}})\bigr{)}\Bigr{)}\\[2.84526pt] &=p_{\ast}^{fake}\Bigl{(}i_{\mathbf{g}}^{!}\,\bigl{(}\mu(u)\bigr{)}^{(\mathbf{g})}\Bigr{)},\end{array}$
@@
It also follows from the definitions of $i^{!}$ and $p_{\ast}$ that for any inclusion $i:\mathcal{X}\to\mathcal{Y}$, each fake push-forward to a point is respected by the wrong-way pull-back $i^{!}$. Thus, push-forwards to points are preserved by $i^{!}$.
@@ 3.6. Compatibility
Given an element $\alpha\in K^{0}(\mathcal{Y})$ and an element $T\in K_{0}(\mathcal{Y})$, we define their pairing by
@@
The following proposition is a consequence of Propositions 3.2 and 3.4.
@@ 4. K-theoretic virtual fundamental cycle
A stably almost complex global Kuranishi chart consists of a quadruple $A=(G,Y,E,s)$, where $G$ is a compact Lie group, $Y$ is a locally $C^{1}$ $G$-manifold with an almost free $G$-action (so that $T^{\ast}_{G}Y$ is stably complex), $E$ is a locally $C^{1}$ $G$-equivariant stably complex vector bundle over $Y$, and $s$ is an equivariant section $Y\to E$.
In [2], Abouzaid, Mclean, and Smith constructed stably almost complex global Kuranishi charts on the compactified moduli space of stable maps ¯$\bar{\mathcal{M}}_{g.n}(X,d)$ to general symplectic targets $X$. We shall make the following modifications and use orbifold Kuranishi charts.
@@
An almost complex global orbifold Kuranishi chart consists of a triple $\mathcal{A}=(\mathcal{Y,E},\sigma)$, where $\mathcal{Y}$ is a smooth, compact orbifold with boundary, $\mathcal{E}$ is a complex smooth orbi-bundle over $\mathcal{Y}$, and $\sigma$ is a continuous section $\mathcal{Y}\to\mathcal{E}$ that vanishes nowhere on $\partial\mathcal{Y}$.
Note that almost complex global orbifold Kuranishi charts exist. Since the real dimensions of $Y$ and the fibers of $E$ have the same parity, we can always add the same $\mathbb{R}^{k}$ to them (and add the identity map $id^{k}:\mathbb{R}^{k}\to\mathbb{R}^{k}$ to the section $s$) to obtain complex $Y$ and $E$. An almost complex global Kuranishi chart $(G,Y,E,s)$ gives an almost complex orbifold Kuranishi chart $(\mathcal{Y,E},\sigma)$ by taking the quotient by the $G$-action and then shrinking to a compact neighborhood of $\sigma^{-1}(0)$. Since $C^{1}$ structures induce smooth structures, we can also modify $Y$ and $E$ to smooth orbifolds and smooth bundles. Note that $s$ is not smooth in general.
By [2] Proposition 4.62 and Lemma 3.4 (see also [1]), global Kuranishi charts $A=(G,Y,E,s)$ for the same moduli space of stable maps of a certain type are connected by a sequence of operations of the following types (or their inverses):
@@
    Free quotients: This operation replaces $(G,Y,E,s)$ with $(G/H,Y/H,E/H,s)$, where $E/H$ is viewed as a $G/H$-bundle over $Y/H$.
@@
    Induced maps: This replaces $(G,Y,E,s)$ with $(G^{\prime},Y\times_{G}G^{\prime},\pi^{\ast}E,\pi^{\ast}s)$, where $\pi:Y\times_{G}G^{\prime}\to Y$ is the projection induced by a slicing $G^{\prime}\to G$.
@@
    Stabilization: Stabilization by a $G$-equivariant vector bundle $\pi:W\to Y$ replaces $(G,Y,E,s)$ with $(G,W,\pi^{\ast}E\oplus\pi^{\ast}W,\pi^{\ast}s\oplus\Delta)$, where $\Delta$ is the tautological “diagonal” section of $W\to\pi^{\ast}W$.
@@
    Open embeddings: This replaces $(G,Y^{\prime},E|_{Y^{\prime}},s|_{Y^{\prime}})$ with $(G,Y,E,s)$ when $Y^{\prime}\subseteq Y$ is an open set containing $s^{-1}(0)$.
In turn, the orbifold Kuranishi charts are related by a sequence of the following operation (and its inverse):
(*) Replace $\mathcal{A}=(\mathcal{Y,E},\sigma)$ by $\mathcal{A}^{\prime}=(\mathcal{W},\mathcal{N}\oplus\mathcal{E},\sigma\oplus\delta)$, where $\pi:\mathcal{N}\to\mathcal{Y}$ is a complex orbi-bundle, $\mathcal{W}\subseteq\mathcal{N}$ is a closed neighborhood of $\sigma^{-1}(0)$, and $\delta$ is the restriction of the diagonal map $\mathcal{N}\to\pi^{\ast}\mathcal{N}$ to $\mathcal{W}$.
4.1. Virtual fundamental cycle of global Kuranishi chart
We define the K-theoretic virtual fundamental cycle $\mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}\in K_{0}(\mathcal{Y})$ of an almost complex global orbifold Kuranishi chart $(\mathcal{Y,E},\sigma)$ as follows. Let $\pi:T^{\ast}\mathcal{Y}\to\mathcal{Y}$ be the projection that forgets the cotangent vector. Let
$\mathcal{S}=\wedge^{\ast}_{\mathbb{C}}T^{\ast}\mathcal{Y},\quad\theta=\wedge^{\ast}_{\mathbb{C}}\mathcal{E},\quad\text{and}\quad\mathcal{O}=\mathcal{S}\otimes\theta.$
Note that $\mathcal{S}$ and $\theta$ are $\mathbb{Z}_{2}$-graded bundles (and hence so is the tensor product $\mathcal{O}$). We add Hermitian structures to $\mathcal{Y}$ and $\mathcal{E}$, and let
@@ 4.1. Virtual fundamental cycle of global Kuranishi chart
Recall that $\varepsilon$ is the operator on $\mathbb{Z}_{2}$-graded objects acting as $(-1)^{a}$ on the direct summand of degree $a$. Also, recall that $\lambda$ is defined as follows: given a section $\tau:\mathcal{X}\to\mathcal{F}$ of a Hermitian bundle, we define $\lambda_{\tau}:\wedge^{\ast}\mathcal{F}\to\wedge^{\ast}\mathcal{F}$ by
$\lambda_{\tau}(\omega)=\tau(x)\wedge\omega+i_{\tau(x)}(\omega)$
for $x\in\mathcal{X}$, where $i_{\tau(x)}$ denotes contraction with $\tau(x)$ via the Hermitian metric on $\mathcal{F}$. Form $\lambda_{\xi}:\pi^{\ast}\mathcal{S}\to\pi^{\ast}\mathcal{S}$ and $\lambda_{\sigma}:\mathcal{E}\to\mathcal{E}$ corresponding to the diagonal map $\xi:T^{\ast}\mathcal{Y}\to\pi^{\ast}T^{\ast}\mathcal{Y}$ and to the section $\sigma:\mathcal{Y}\to\mathcal{E}$. Observe that $\lambda_{\xi}$ (respectively, $\lambda_{\sigma}$) is of degree $1$, so $p$ is also of degree $1$. Moreover, $\lambda_{\xi}$ and $\lambda_{\sigma}$ are self-dual and square to $|\xi|^{2}\cdot 1$ and $|\sigma(x)|^{2}\cdot 1$, respectively; hence,
$p^{\ast}p=(|\xi|^{2}+|\sigma(x)|^{2})\cdot 1.$
Thus, $p$ is invertible outside the compact set
=$\mathcal{M}=\{x\in\mathcal{Y}:|\sigma(x)|=0\text{ and }|\xi|=0\},$
and therefore represents an element in $K^{0}(T^{\ast}\mathcal{Y})$.
It follows from Proposition 3.3 that, in terms of the K-theoretic virtual fundamental cycle, the operation (*) corresponds to first pushing forward by $i_{\ast}$ (defined in Section 2.3) and then pulling back by $j^{!}$ (defined in Section 3.3).
@@
(For Kuranishi charts) If $\sigma$ is transversal to the zero section 00, then
$\mathcal{O}^{vir}_{\mathcal{Y,E},\sigma}=i_{\ast}[\sigma^{-1}(0)].$
@@
Note that the space $\mathcal{M}=\sigma^{-1}(0)$ is stably almost complex. We can define its K-theoretic fundamental cycle by Bott periodicity, since $K_{0}(\Sigma^{k}\mathcal{M})=K_{0}(\mathcal{M})$. This identification is exactly our $i^{!}$ for the inclusion $i:\mathcal{M}\to B(\mathcal{M}\times\mathbb{R}^{2k})$. Next, by part (2) of Proposition 3.3, the push-forwards $i_{\ast}$ and the wrong-way pull-back $i^{!}$ are intertwined by the choose-an-operator map, thus proving the theorem. ∎
@@
(For Kuranishi charts)
$\langle\mathcal{O}^{vir}_{\sigma},\mathcal{V}\rangle=\operatorname{ind}[\mathcal{E}\otimes\pi^{\ast}\mathcal{V}]_{+}.$
4.2. Virtual fundamental cycle of virtual orbifolds
We further define the virtual fundamental cycle as
$i^{!}(\mathcal{O}^{vir}_{\sigma})\in K_{0}(\mathcal{M}).$
Proposition 4.1.
This virtual fundamental cycle is independent of the global Kuranishi chart.
Proof.
This follows immediately from the fact that $i^{!}j^{!}=(ji)^{!}$. ∎
We define $p_{\ast}$ by the same formula as in Definition 3.1, with the “normal bundles” replaced by
$\mathcal{N}_{\mathbf{g}}^{\mathcal{M}}=\mathcal{N}_{\mathbf{g}}\ominus\mathcal{E}_{\mathbf{g}}|_{\mathcal{M}^{\mathbf{g}}}$
in an arbitrary global chart. Then we immediately have the following proposition.
Proposition 4.2.
$\mathcal{N}_{\mathbf{g}}^{\mathcal{M}}$ is independent of the global Kuranishi chart $(\mathcal{Y,E},\sigma)$.
We also have
Theorem 2.
(For orbispaces with almost complex global Kuranishi chart) Let $i:\mathcal{M}\to\mathcal{Y}$ be the inclusion. Then
$\langle\mathcal{O}^{vir}_{\sigma},V\rangle=\langle\mathcal{O}^{vir}_{\mathcal{M}},i^{\ast}V\rangle.$
Proof.
We only need to prove that the index $p_{\ast}^{fake}$ on each $K$-homology element $e_{\mathbf{g}}^{-1}i_{\mathbf{g}}^{!}\varphi^{(g)}$ coincides. Note that this index depends only on the $\mathbb{Z}_{2}$-graded Hilbert space $H$ and the operator $T$ involved, and is independent of the module structure $\phi$ over the ring of continuous functions. To obtain the same (virtual) Hilbert spaces and Fredholm operators, for each $\mathbf{g}$ we can consider
$\frac{e_{\mathbf{g}}^{-1}i_{\mathbf{g}}^{!}\varphi^{(g)}}{\wedge^{\ast}(\mathcal{E}|_{\mathcal{Y}^{\mathbf{g}}}/\mathcal{E}^{\mathbf{g}})^{\ast}}$
over $\mathcal{Y}^{\mathbf{g}}$ and $\mathcal{E}|_{\mathcal{Y}^{\mathbf{g}}}/\mathcal{E}^{\mathbf{g}}$. The coincidence of indices now follows from the end of Section 3.5. ∎